|
Group Selection Page
HELP
Controls Group
ITDM
Vacuum
WP78
SPB
SQS
XO
ToBeMigrated2XWiki
ToBeMigrated2GitlabPages
ToBeMigratedByGroup
Migrated2Zulip
Archived2PDF
not2MigrateOrArchive
|
|
SCS
SXP Migrated2Zulip
Photon Commissioning Team
XRO
DOC
HED Migrated2Zulip
WP74
MID
PSPO_XO
XO_Operation
SQS Migrated2Zulip
FXE Migrated2Zulip
DET
Sample Environment
SPB Migrated2Zulip
|
|
SASE1 Status and SASE3 Readiness
Technical integration and coordination
|
| Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
|
Message ID: 74
Entry time: 31 Jul 2019, 10:46
|
| Author: |
Adriano Violante |
| Location: |
General |
| Severity: |
Information |
| Subject: |
Minutes of the Technical integration meeting 31.07.19 |
|
|
XO
FXE
SPB/SFX
- Loop 12 has to be updated ==> Done.
- SASE1 M3 issues with the pitch motor lead to FXE losing beamtime. EEE to follow it up. ==> Basically fixed. Some remaining issues will be fixed during the shutdown week in August.
- Some issues reading the encoders of the JJ XRay CSLIT. EEE checks whether similar issues are also present in other instruments.
SCS
SQS
- Commissioning of smaract stages ongoing.
MID
- By the end of the week MID should send the requirements for the vacuum interlock to EEE. ==> tests are ongoing. AS will organize a meeting with CAS, VAC and EEE to define a test procedure to be followed by the instruments when testing the vacuum interlocks. ==> to be done.
HED
- New update of Loop 6 will be done propably today after new interlock definitions have been delivered.
- Hexapod: Loop 9 has been updated. Loop 7 to follow.
Optical lasers
VAC
XRO
XPD
DET
EEE
- Crash of PLCs because of not a number values. The fix on the PLC side requires the update of the loops framework and can be done only in the August shutdown. By the end of the week a filtering tool will be availbale in Karabo.==> Karabo client has been deployed in the tunnels and FXE. The other instruments will get it during the August shutdown.
- XPD reported some issues with digitizers for MCPs. It should be fixed already by now.
CAS
- CAS received host alerts regarding sa2-br-sys-con-1, sa2-br-sys-con-2, sa2-br-sys-con-5 resources usage. To avoid the worst issues during experiments, we will do maintenance on these machines today from 12:00 to 14:00.
- CAS/SQS working on karabo scenes to display the LaserMet laser isnterlock status.
ITDM
SRP
TS
AoB
- RBean reported that by the end of the week a macro will be run by HSinn to decouple machine interlocks in SASE1 and SASE3. For more details AGaller should be addressed.
|